OpenStudy (anonymous):

~Help Please~ Write an equasion in slope intercept form for the line that passes throuh the point (-4,6) and is Perpendicular to y=8x-7 Show work PLEASE <3

OpenStudy (anonymous):

Perpendicular to y=8x-7 means you have to take slope -1/8 y-6=-1/8(x+4) y=-1/8x +(6-1/2)=-1/8x +11/2

OpenStudy (anonymous):

last one answer?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

y=-1/8x+11/2

OpenStudy (anonymous):

thank you